A total of 1135 people attended a concert. Floor tickets cost $40 each, while balcony tickets cost $28 each. If a total of $40,480 was collected, how many of each type of ticket were sold?
Answer by checkley77(12844)

F+B=1,135 OR F=1,135-B
40F+28B=40,480
40(1,135-B)+28B=40,480
45,400-40B+28B=40,480
-12B=40,480-45,400
-12B=-4,920
B=-4,920/-12
B=410 NUMBER OF BALCONY TICKETS WERE SOLD.
1,135-410=725 NUMBER OF FLOOR TICKETS WERE SOLD.
40*725+28*410=40,480
29,000+11,480=40,480
40,480=40,480